OpenStudy (anonymous):

you roll two dice. What is the probability of rolling a 3 on the first die and a 5 on the second one?

OpenStudy (anonymous):

(1-5/6)(1-5/6)

OpenStudy (anonymous):

event A = rolling die one and probability that you donot get 3 is P(A)= 5/6, so the prob of getting 3 is 1-p(A)=1/6 event B = rolling die two and prob that you donot get 5 is P(b)=5/6, so the prob of gtting a 5 is 1-p(b)=1/6 The total probability is P(a)P(b) BECOZ these are INDEPENDENT EVENTS = 1/6*1/6=1/36